# Non-local finite-depth circuits for constructing SPT states and quantum cellular automata

### 04 Dec 2023 v3

@@ Abstract
The exploration of topological phases of matter has been a major theme of modern condensed matter physics (for introductions, see Wen (2004, 2017)), with far-reaching implications for quantum information and quantum computation (for introductions, see Bombín (2013); Lahtinen and Pachos (2017)). These phases of matter, when defined on lattices, have been classified using the complexity of preparation of the associated ground states using local quantum circuits Chen et al. (2010a); a state that can be accessed by starting with a product state and acting with a finite-depth quantum circuit of geometrically local gates (FDQC) can be said to be ‘easy’ to prepare, and one that cannot be, is difficult. This characterization is intimately related to the notion of topological stability, which states that topological phases are robust to geometrically local perturbations (which should be symmetry restricted in the case of symmetry-protected topological (SPT) phases Senthil (2015); Chen et al. (2013)), since topological stability implies that topologically non-trivial states cannot be connected to trivial states by a (symmetric) FDQC Bravyi et al. (2006); Chen et al. (2010a). Recently, the notion of FDQCs in relation to topological phases has been extended in various directions, such as allowing projective measurements Tantivasadakarn et al. (2022a); Piroli et al. (2021a); Bravyi et al. (2022); Tantivasadakarn et al. (2022b), or by extending to linear depth quantum circuits Huang and Chen (2015); Liu et al. (2022); Wang et al. (2022), but still demanding geometric locality. The geometric locality is, of course, a very natural constraint to impose in traditional solid-state settings. However, we are witnessing the rapid development of experimental capabilities in the fields of quantum simulation and noisy intermediate-scale quantum devices (for a recent review, see Altman et al. (2021)), which have come to provide a new context in which to explore topological phases (see e.g. Semeghini et al. (2021a); Marques et al. (2022)). In this new setting, the geometric locality is not necessarily guaranteed and there can arise $k$-local interactions i.e., interactions that are few-body (acting on no more than $k$ degrees of freedom), but not restricted to be geometrically local (see e.g. Fowler and Martinis (2014); Richerme et al. (2014); Vaidya et al. (2018); Bourassa et al. (2021)). Do the results on the difficulty of topological state preparation survive in this new setting, and related, does the classification of topological phases of matter and the associated notion of topological stability survive in a setting where there can be $k$-local perturbations?
@@
Another topological classification that may be affected by $k$-local interactions is that of locality-preserving unitary operators, also known as quantum cellular automata (QCA) Farrelly (2020a). An FDQC is an obvious example of a QCA, but there are also QCA that cannot be written as an FDQC, such as the lattice translation operator. The topological classification of phases of QCA arises from the question of whether two QCA can be smoothly connected along a locality-preserving path, with the topologically trivial phase consisting of FDQCs as they can be connected to the identity. In one and two dimensions, this classification is given by an index theory that essentially shows that every QCA can be decomposed as an FDQC and a translation Gross et al. (2012); Freedman and Hastings (2020). In three dimensions, the classification is incomplete as there appear to be QCA that are neither FDQCs nor translations, but significant progress is being made Haah et al. (2022); Haah (2021); Shirley et al. (2022). The classification of QCA also plays a role in the classification of topological phases Stephen et al. (2019); Haah (2021); Haah et al. (2022); Shirley et al. (2022); Fidkowski et al. (2020), particularly Floquet phases Po et al. (2016); Fidkowski et al. (2019); Po et al. (2017).
@@
In this work, we examine the fate of the ‘topological stability’ of topological phases and QCA in the presence of $k$-local perturbations by asking whether one can construct a reference state or QCA in the putatively topological phase using an FDQC made out of $k$-local (but not geometrically local) gates. We show that SPT phases with global or subsystem symmetries are not stable to $k$-local perturbations by explicitly constructing FDQCs of symmetric $k$-local gates that trivialize the fixed-point states by mapping them to symmetric product states. Since all states within a given phase can be connected to the fixed-point state by a symmetric FDQC Chen et al. (2010b), our circuits for fixed-point states imply the existence of similar circuits that disentangle any state in a given SPT phase. We also argue that SPT order should be unstable to generic symmetric $k$-local noise. Similarly, we show that every translationally-invariant QCA on a periodic lattice in any spatial dimension can be realized as an FDQC of $k$-local gates that commute with the same symmetries as the QCA. Therefore, all QCA belong to the same trivial phase when $k$-local gates are allowed. However, $k$-local perturbations do not trivialize everything - indeed it is known that topological order (such as the toric code model) is stable to $k$-local perturbations Aharonov and Touati (2018). More generally, it is known that code states of quantum error correcting codes can not be disentangled in finite-depth by $k$-local unitary gates Bravyi et al. (2006); Aharonov and Touati (2018) (see Appendix D for a brief review of the proof). We argue that fracton order and Floquet topological codes are also stable to $k$-local perturbations, and we conjecture the same is true of SPT order protected by higher-form symmetries.
@@ II.4 SPT phases with higher-form symmetries
We first observe that the folding argument used for global symmetries does not carry over to the case of higher-form symmetries. This is because folding in this case does not correspond to stacking in the usual sense. Recall that stacking requires the symmetry to act in the same way on the two layers. When we fold a system with global symmetry, acting with the symmetry on the whole system automatically has identical action on the two layers of the fold. However, this is not the case when we fold a system with higher-form symmetry, since the symmetry generators, being local, can act independently on either layer. Because of this, the bulk of the folded system resembles two stacked SPT states, each with their own independent symmetry, and such a system has non-trivial SPT order.
Second, we observe that any $k$-local interaction which commutes with a higher-form symmetry is also locally symmetric, meaning the gates can be decomposed into a sum of tensor products of geometrically local symmetric unitaries. This is due to the simple fact that the higher-form symmetry group itself contains operators which act non-trivially only in local regions of space. This is much different from the case of global symmetries, where a symmetric $k$-local interaction can violate symmetry locally while still preserving it globally. Because of this, the class of $k$-local gates which we can employ is severely restricted compared to the case of global symmetry. Indeed, the $\mathrm{FDQC}_{k}$’s used to disentangle the SPT states above only commute with the symmetry globally as they contain operators such as a long range $ZZ$ pair which can transfer symmetry charge over long distances. In Sec. IV, we argue that the fact that symmetric $k$-local interactions can violate the symmetry in local regions of space is crucial to their ability to destroy SPT order. Accordingly, we expect that $k$-local gates which preserve the symmetry locally, which is always the case for higher-form symmetric gates, are not sufficient to destroy SPT order.
Finally, examining the boundary of a higher-form SPT also suggests that it may be $k$-local non-trivial. Higher-form SPTs in 3D, for example, can support topologically ordered boundary theories such as a 2D toric code appearing on the boundary of the 3D cluster state Walker and Wang (2012); Yoshida (2016); Roberts and Bartlett (2020). However, unlike in the boundaries of systems with global symmetry discussed above, $k$-local interactions cannot trivialize this boundary theory because the 2D toric code, and any number of stacks of it, is $k$-local non-trivial Aharonov and Touati (2018).
@@ III Unified construction of symmetric $k$-local circuits for QCA
In this section, we show that locality-preserving unitary operators, also known as quantum cellular automata (QCA) Farrelly (2020b), can be implemented as $\mathrm{FDQC}_{k}$’s. We give a generic construction of these $\mathrm{FDQC}_{k}$’s that works in any spatial dimension $d$. We only deal with translationally-invariant QCA, but this condition can likely be relaxed to a certain extent. We also assume the QCA acts on a lattice with mirror symmetry in all $d$ directions.
@@
An FDQC is a trivial examples of a QCA. However, there are also QCA that cannot be expressed as an FDQC, such as the 1D shift QCA $Q_{S}$ which acts on any operator $O_{i}$ supported on site $i$ by translating it by one site, $Q_{S}O_{i}Q_{S}^{-1}=O_{i+1}$. When only local gates are employed, implementing this operation requires QC whose depth grows linearly with system size. In 1D and 2D, it has been shown that all QCA are composed of shifts and FDQCs Gross et al. (2012); Freedman and Hastings (2020). In 3D, however, there are believed to be QCA which are neither shifts nor circuits Haah et al. (2022); Haah (2021); Shirley et al. (2022). Our result implies that all of these QCA become trivial in the $k$-local scenario, i.e., they can be written as $\mathrm{FDQC}_{k}$’s. Therefore, the topological classification of QCA collapses to a single trivial phase when $k$-local gates are allowed. Additionally, if the QCA commutes with some global symmetry, it may not be possible to write it as a symmetric FDQC, even if it is an FDQC. Nonetheless, the $\mathrm{FDQC}_{k}$ we construct consists of symmetric gates (in the case of global symmetries), so this result gives an alternative demonstration of the triviality of SPT order by applying our construction to the QCA which generate the SPT states, such as those in Eq. 2 and Eq. 5.
@@
The main ideas which lead to this result are the following. Given a QCA $Q$ acting in $d$ spatial dimensions, we start with the known fact that $Q\otimes Q^{-1}$ acting on two copies of a system can be realized by an FDQC Arrighi et al. (2011); Farrelly (2020b). By truncating this FDQC to a finite region of space, we obtain a $k$-local circuit acting on a single conjoined system. On one half of this system, the circuit acts like $Q$, while on the other half, it acts like $Q^{-1}$. We then show that $Q^{-1}$ is equivalent to the spatial inversion of $Q$ in one direction, which we denote as $\bar{Q}$, up to a circuit of $(d-1)$-dimensional QCA that can be applied in two parallel layers. The action of $Q$ on one half and $\bar{Q}$ on the other is nothing more than $Q$ applied to a periodic system. An inductive argument then concludes that $Q$ can be implemented as a $\mathrm{FDQC}_{k}$. This idea is demonstrated in Fig. 4 for the simple example of the shift QCA. In this case, the second part of the argument is not needed as $\bar{Q}$ is already equal to $Q^{-1}$. We work out three explicit examples of the general construction, including the shift QCA, in Appendix B.
@@ III.1 Construction from 1D Margolus representation
Consider a QCA $Q$ which acts on a Hilbert space $\mathcal{H}=(\mathbb{C}^{d})^{\otimes N}$. Now construct a doubled Hilbert space $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ made of two copies of $\mathcal{H}$. For every degree of freedom $i$ in $\mathcal{H}$, we have two degrees of freedom $[i]_{A}$ and $[i]_{B}$ in $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$. Let $Q_{A}$ ($Q_{B}$) denote $Q$ acting on $\mathcal{H}_{A}$ ($\mathcal{H}_{B}$) and write $V=Q_{B}^{-1}Q_{A}$. Now write $\mathcal{S}_{AB}=\prod_{i}S_{i}$ where $S_{i}$ is the SWAP operation that exchanges sites $[i]_{A}$ and $[i]_{B}$. For now, we imagine that $A$ and $B$ are geometrically close such that $S_{i}$ is a local operator. Then we have Arrighi et al. (2011); Farrelly (2020b),
@@
$V_{i}=Q_{A}S_{i}Q_{A}^{-1}.$ (11)
Since $S_{i}$ is a local operator and $Q$ is locality-preserving, $V_{i}$ is a local operator. Since the $V_{i}$ all commute with each other for all $i$, the above formula can be parallelized into a finite-depth circuit realizing $V$.
@@
Figure 5: The truncated circuit $V_{R}$ where the 1D QCA $Q$ is expressed in standard form. The downwards- (upwards-) shifted vertical lines correspond to degrees of freedom in the $A$ ($B$) subsystem. The lighter-colored rectangles represent the inverses $u^{-1}$ and $v^{-1}$ as indicated in the first expression. An arrow between two lines indicates the SWAP operation $S_{i}$ between degrees of freedom in the two subsystems. In the second expression, most gates have canceled pairwise, leaving only those shown. For visual clarity, the vertical lines have been shortened. The third expression is equivalent to the second expression up to the symmetric finite-depth quantum circuits (FDQCs) $W_{1}$ and $W_{2}$; see Eq. III.1. Note that the strict equivalence up to symmetric FDQCs is in one spatial dimension, $d=1$. In higher dimensions ($d>1$), we use the compactified picture such that the equivalence is up to a circuit that is finite-depth along the non-compactified direction and in general, a $(d-1)$-dimensional QCA along the $d-1$ compactified directions; see main text.
@@
This is depicted in the first two diagrams in Fig. 5. As is clear from the figure, these unitaries can be parallelized into two layers of disjoint operators. Furthermore, we see that the degrees of freedom within $R$ have been stitched into a single periodic 1D system, a ring of length $L=2(d-c+1)+2$ with spins in the $A$ ($B$) subsystem forming the “front” (“back”) of the ring that we denote as $\mathcal{R}$. Whenever we use the notation $\mathcal{R}$, we imagine it as representing a finite periodic array of sites that can be ordered counter-clockwise as $[c]_{A},\dots,[d+1]_{A},[d]_{B},[d-1]_{B},\dots,[c]_{B},[c-1]_{A}$.
The operator $V_{R}$ is not exactly $Q$ acting on the finite periodic system defined by $\mathcal{R}$, which is depicted in the third expression in Fig. 5. Instead, $V_{R}$ realizes $Q$ on the front of $\mathcal{R}$ and $\bar{Q}^{-1}$ on the back of $\mathcal{R}$, with the two operators being blended near the edges. To fix this, we show that $V_{R}$ is equivalent to $Q$ up to composition with an FDQC. Given any QCA $Q$, we show that its inverse $Q^{-1}$ is related to its orientation-reversed self $\bar{Q}$ by an FDQC. This is intuitively clear for 1D QCA such as the shift QCA: reversing the direction of the shift is the same as inverting it. Generally, $\bar{Q}$ will not equal $Q^{-1}$, but they will differ only by an FDQC in 1D  [^2] Indeed, if we let $\operatorname{ind}(Q)$ be the rational-valued index of $Q$ as defined in Ref. Gross et al. (2012), then it is straightforward to show that $\operatorname{ind}(Q^{-1})=\operatorname{ind}(Q)^{-1}=\operatorname{ind}(\bar{Q})$, which implies that $Q^{-1}$ and $\bar{Q}$ differ by an FDQC in 1D Gross et al. (2012).
To explicitly construct the circuit which maps $\bar{Q}$ to $Q^{-1}$, we introduce the unitary $w=\bar{v}u$ where $\bar{v}$ is the spatial reversal (opposite orientation) of $v$ obtained by exchanging the left and right input and output Hilbert spaces. Graphically,
@@
Thus $Q^{-1}$ and $\bar{Q}$ are related by composition with FDQCs. Using this, we define the circuits acting on the “back” of $\mathcal{R}$,

### 11 Jan 2024 v4

@@ Abstract
The exploration of topological phases of matter has been a major theme of modern condensed matter physics (for introductions, see Wen (2004, 2017)), with far-reaching implications for quantum information and quantum computation (for introductions, see Bombín (2013); Lahtinen and Pachos (2017)). These phases of matter, when defined on lattices, have been classified using the complexity of preparation of the associated ground states using local quantum circuits Chen et al. (2010); a state that can be accessed by starting with a product state and acting with a finite-depth quantum circuit of geometrically local gates (FDQC) can be said to be ‘easy’ to prepare, and one that cannot be, is difficult. This characterization is intimately related to the notion of topological stability, which states that topological phases are robust to geometrically local perturbations (which should be symmetry restricted in the case of symmetry-protected topological (SPT) phases Senthil (2015); Chen et al. (2013)), since topological stability implies that topologically non-trivial states cannot be connected to trivial states by a (symmetric) FDQC Bravyi et al. (2006); Chen et al. (2010). Recently, the notion of FDQCs in relation to topological phases has been extended in various directions, such as allowing projective measurements Piroli et al. (2021a); Bravyi et al. (2022); Tantivasadakarn et al. (2022), or by extending to linear depth quantum circuits Huang and Chen (2015); Liu et al. (2022); Wang et al. (2022), but still demanding geometric locality. The geometric locality is, of course, a very natural constraint to impose in traditional solid-state settings. However, we are witnessing the rapid development of experimental capabilities in the fields of quantum simulation and noisy intermediate-scale quantum devices (for a recent review, see Altman et al. (2021)), which have come to provide a new context in which to explore topological phases (see e.g. Semeghini et al. (2021a); Marques et al. (2022)). In this new setting, the geometric locality is not necessarily guaranteed and there can arise $k$-local interactions i.e., interactions that are few-body (acting on no more than $k$ degrees of freedom), but not restricted to be geometrically local (see e.g. Fowler and Martinis (2014); Richerme et al. (2014); Vaidya et al. (2018); Bourassa et al. (2021)). Do the results on the difficulty of topological state preparation survive in this new setting, and related, does the classification of topological phases of matter and the associated notion of topological stability survive in a setting where there can be $k$-local perturbations?
@@
Another topological classification that may be affected by $k$-local interactions is that of locality-preserving unitary operators, also known as quantum cellular automata (QCA) Farrelly (2020). An FDQC is an obvious example of a QCA, but there are also QCA that cannot be written as an FDQC, such as the lattice translation operator. The topological classification of phases of QCA arises from the question of whether two QCA can be smoothly connected along a locality-preserving path, with the topologically trivial phase consisting of FDQCs as they can be connected to the identity. In one and two dimensions, this classification is given by an index theory that essentially shows that every QCA can be decomposed as an FDQC and a translation Gross et al. (2012); Freedman and Hastings (2020). In three dimensions, the classification is incomplete as there appear to be QCA that are neither FDQCs nor translations, but significant progress is being made Haah et al. (2022); Haah (2021); Shirley et al. (2022). The classification of QCA also plays a role in the classification of topological phases Stephen et al. (2019); Haah (2021); Haah et al. (2022); Shirley et al. (2022); Fidkowski et al. (2020), particularly Floquet phases Po et al. (2016); Fidkowski et al. (2019); Po et al. (2017).
@@
In this work, we examine the fate of the ‘topological stability’ of topological phases and QCA in the presence of $k$-local perturbations by asking whether one can construct a reference state or QCA in the putatively topological phase using an FDQC made out of $k$-local (but not geometrically local) gates. We show that SPT phases with global or subsystem symmetries are not stable to $k$-local perturbations by explicitly constructing FDQCs of symmetric $k$-local gates that trivialize the fixed-point states by mapping them to symmetric product states. Since all states within a given phase can be connected to the fixed-point state by a symmetric FDQC Chen et al. (2010), our circuits for fixed-point states imply the existence of similar circuits that disentangle any state in a given SPT phase. We also argue that SPT order should be unstable to generic symmetric $k$-local noise. Similarly, we show that every translationally-invariant QCA on a periodic lattice in any spatial dimension can be realized as an FDQC of $k$-local gates that commute with the same symmetries as the QCA. Therefore, all QCA belong to the same trivial phase when $k$-local gates are allowed. However, $k$-local perturbations do not trivialize everything - indeed it is known that topological order (such as the toric code model) is stable to $k$-local perturbations Aharonov and Touati (2018). More generally, it is known that code states of quantum error correcting codes can not be disentangled in finite-depth by $k$-local unitary gates Bravyi et al. (2006); Aharonov and Touati (2018) (see Appendix D for a brief review of the proof). We argue that fracton order and Floquet topological codes are also stable to $k$-local perturbations, and we conjecture the same is true of SPT order protected by higher-form symmetries.
@@ II.4 SPT phases with higher-form symmetries
We first observe that the folding argument used for global symmetries does not carry over to the case of higher-form symmetries. This is because folding, in this case, does not correspond to stacking in the usual sense. Recall that stacking requires the symmetry to act in the same way on the two layers. When we fold a system with global symmetry, acting with the symmetry on the whole system automatically has identical action on the two layers of the fold. However, this is not the case when we fold a system with higher-form symmetry since the symmetry generators, being local, can act independently on either layer. Because of this, the bulk of the folded system resembles two stacked SPT states, each with its own independent symmetry, and such a system has a non-trivial SPT order.
Second, we observe that any $k$-local interaction that commutes with a higher-form symmetry is also locally symmetric, meaning the gates can be decomposed into a sum of tensor products of geometrically local symmetric unitaries. This is due to the simple fact that the higher-form symmetry group itself contains operators that act non-trivially only in local regions of space. This is much different from the case of global symmetries, where a symmetric $k$-local interaction can violate symmetry locally while still preserving it globally. Because of this, the class of $k$-local gates that we can employ is severely restricted compared to the case of global symmetry. Indeed, the $\mathrm{FDQC}_{k}$’s used to disentangle the SPT states above only commute with the symmetry globally as they contain operators such as a long-range $ZZ$ pair, which can transfer symmetry charge over long distances. In Sec. IV, we argue that the fact that symmetric $k$-local interactions can violate the symmetry in local regions of space is crucial to their ability to destroy SPT order. Accordingly, we expect that $k$-local gates that preserve the symmetry locally, which is always the case for higher-form symmetric gates, are insufficient to destroy SPT order.
Finally, examining the boundary of a higher-form SPT also suggests that it may be $k$-local non-trivial. Higher-form SPTs in 3D, for example, can support topologically ordered boundary theories such as a 2D toric code appearing on the boundary of the 3D cluster state Walker and Wang (2012); Yoshida (2016); Roberts and Bartlett (2020). However, unlike in the boundaries of systems with global symmetry discussed above, $k$-local interactions cannot trivialize this boundary theory because the 2D toric code, and any number of stacks of it, are $k$-local non-trivial Aharonov and Touati (2018).
@@ III Unified construction of symmetric $k$-local circuits for QCA
In this section, we show that locality-preserving unitary operators, also known as quantum cellular automata (QCA) Farrelly (2020), can be implemented as $\mathrm{FDQC}_{k}$’s. We give a generic construction of these $\mathrm{FDQC}_{k}$’s that works in any spatial dimension $d$. We only deal with translationally-invariant QCA, but this condition can likely be relaxed to a certain extent. We also assume the QCA acts on a lattice with mirror symmetry in all $d$ directions.
@@
An FDQC is a trivial example of a QCA. However, there are also QCA that cannot be expressed as an FDQC, such as the 1D shift QCA $Q_{S}$ which acts on any operator $O_{i}$ supported on site $i$ by translating it by one site, $Q_{S}O_{i}Q_{S}^{-1}=O_{i+1}$. When only local gates are employed, implementing this operation requires QC whose depth grows linearly with system size. In 1D and 2D, it has been shown that all QCA are composed of shifts and FDQCs Gross et al. (2012); Freedman and Hastings (2020). In 3D, however, there are believed to be non-trivial QCA which are neither shifts nor circuits Haah et al. (2022); Haah (2021); Shirley et al. (2022). Our result constructs $\mathrm{FDQC}_{k}$’s for all QCA, which, to the best of our knowledge, provides the first circuit representation of nontrivial 3D QCA. The existence of these $\mathrm{FDQC}_{k}$’s implies that all QCA become trivial in the $k$-local scenario, such that the topological classification of QCA collapses to a single trivial phase when $k$-local gates are allowed. Additionally, if the QCA commutes with some global symmetry, it may not be possible to write it as a symmetric FDQC, even if it is an FDQC. Nonetheless, the $\mathrm{FDQC}_{k}$ we construct consists of symmetric gates (in the case of global symmetries), so this result gives an alternative demonstration of the triviality of SPT order by applying our construction to the QCA which generate the SPT states, such as those in Eq. 2 and Eq. 5.
@@
The main ideas which lead to this result are the following. Given a QCA $Q$ acting in $d$ spatial dimensions, we start with the known fact that $Q\otimes Q^{-1}$ acting on two copies of a system can be realized by an FDQC Arrighi et al. (2011); Farrelly (2020). By truncating this FDQC to a finite region of space, we obtain a $k$-local circuit acting on a single conjoined system. On one half of this system, the circuit acts like $Q$, while on the other half, it acts like $Q^{-1}$. We then show that $Q^{-1}$ is equivalent to the spatial inversion of $Q$ in one direction, which we denote as $\bar{Q}$, up to a circuit of $(d-1)$-dimensional QCA that can be applied in two parallel layers. The action of $Q$ on one half and $\bar{Q}$ on the other is nothing more than $Q$ applied to a periodic system. An inductive argument then concludes that $Q$ can be implemented as a $\mathrm{FDQC}_{k}$. This idea is demonstrated in Fig. 4 for the simple example of the shift QCA. In this case, the second part of the argument is not needed as $\bar{Q}$ is already equal to $Q^{-1}$. We work out three explicit examples of the general construction, including the shift QCA, in Appendix B.
@@ III.1 Construction from 1D Margolus representation
Consider a QCA $Q$ which acts on a Hilbert space $\mathcal{H}=(\mathbb{C}^{d})^{\otimes N}$. Now construct a doubled Hilbert space $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ made of two copies of $\mathcal{H}$. For every degree of freedom $i$ in $\mathcal{H}$, we have two degrees of freedom $[i]_{A}$ and $[i]_{B}$ in $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$. Let $Q_{A}$ ($Q_{B}$) denote $Q$ acting on $\mathcal{H}_{A}$ ($\mathcal{H}_{B}$) and write $V=Q_{B}^{-1}Q_{A}$. Now write $\mathcal{S}_{AB}=\prod_{i}S_{i}$ where $S_{i}$ is the SWAP operation that exchanges sites $[i]_{A}$ and $[i]_{B}$. For now, we imagine that $A$ and $B$ are geometrically close such that $S_{i}$ is a local operator. Then we have Arrighi et al. (2011); Farrelly (2020),
@@
$V_{i}=Q_{A}^{-1}S_{i}Q_{A}.$ (11)
Since $S_{i}$ is a local operator, and $Q_{A}$ is locality-preserving, $V_{i}$ is a local operator. Since the $V_{i}$ all commute with each other for all $i$, the above formula can be parallelized into a finite-depth circuit realizing $V$.
@@
Figure 5: The truncated circuit $V_{R}$ where the 1D QCA $Q$ is expressed in standard form. The downwards- (upwards-) shifted vertical lines correspond to degrees of freedom in the $A$ ($B$) subsystem. The lighter-colored rectangles represent the inverses $u^{-1}$ and $v^{-1}$ as indicated in the first expression. An arrow between two lines indicates the SWAP operation $S_{i}$ between degrees of freedom in the two subsystems. In the second expression, most gates have canceled pairwise, leaving only those shown. The degrees of freedom acted on by the remaining gates form the ring $\mathcal{R}$. For visual clarity, the vertical lines have been shortened. The third expression is equivalent to the second expression up to the symmetric finite-depth quantum circuits (FDQCs) $W_{1}$ and $W_{2}$; see Eq. III.1. Note that the strict equivalence up to symmetric FDQCs is in one spatial dimension, $d=1$. In higher dimensions ($d>1$), we use the compactified picture such that the equivalence is up to a circuit that is finite-depth along the non-compactified direction and in general, a $(d-1)$-dimensional QCA along the $d-1$ compactified directions; see main text.
@@
$=$ $\left(\prod_{i=c}^{d}S_{i}\right)Q_{A}^{-1}\left(\prod_{i=c}^{d}S_{i}\right)Q_{A}$
@@
The second and third equalities above are depicted on the left and middle of Fig. 5, respectively. As is clear from Fig. 5, the $u$, $u^{-1}$, $v$, $v^{-1}$ operators can be parallelized into two layers of disjoint operators. Namely, the $u$ and $v^{-}1$ operators can act in parallel in the first layer, and the $v$ and $u^{-1}$ operators act in the second layer. Furthermore, we see that the degrees of freedom within the range $[c-1,d+1]$ have been stitched into a single periodic 1D system, a ring of length $L=2(d-c+1)+2$ with spins in the $A$ ($B$) subsystem forming the “front” (“back”) of the ring that we denote as $\mathcal{R}$. Whenever we use the notation $\mathcal{R}$, we imagine it as representing a finite periodic array of sites that can be ordered counter-clockwise as $[c]_{A},\dots,[d+1]_{A},[d]_{B},[d-1]_{B},\dots,[c]_{B},[c-1]_{A}$.
The operator $V_{R}$ is not exactly $Q$ acting on the finite periodic system defined by $\mathcal{R}$, which is depicted in the third expression in Fig. 5. Instead, $V_{R}$ realizes $Q$ on the front of $\mathcal{R}$ and $\bar{Q}^{-1}$ on the back of $\mathcal{R}$, with the two operators being blended near the edges. To fix this, we show that $V_{R}$ is equivalent to $Q$ up to composition with an FDQC. Given any 1D QCA $Q$, we show that its inverse $Q^{-1}$ is related to its orientation-reversed self $\bar{Q}$ by an FDQC. This is intuitively clear for 1D QCA such as the shift QCA: reversing the direction of the shift is the same as inverting it. Generally, $\bar{Q}$ will not equal $Q^{-1}$, but they will differ only by an FDQC in 1D  [^2] Indeed, if we let $\operatorname{ind}(Q)$ be the rational-valued index of $Q$ as defined in Ref. Gross et al. (2012), then it is straightforward to show that $\operatorname{ind}(Q^{-1})=\operatorname{ind}(Q)^{-1}=\operatorname{ind}(\bar{Q})$, which implies that $Q^{-1}$ and $\bar{Q}$ differ by an FDQC in 1D Gross et al. (2012)..
To explicitly construct the circuit which maps $\bar{Q}$ to $Q^{-1}$, we introduce the unitary $w=\bar{v}u$ where $\bar{v}$ is the spatial reversal (opposite orientation) of $v$ obtained by exchanging the left and right input and output Hilbert spaces. That is, $v_{i,j}=\bar{v}_{j,i}$. Graphically,
@@
Since the $w$ ($\bar{w}$) gates on either side of $Q^{-1}$ are non-overlapping, they can be applied in parallel. Therefore, $Q^{-1}$ and $\bar{Q}$ are related by composition with FDQCs. Using this, we define the circuits acting on the “back” of $\mathcal{R}$,